a.) Find the area of the triangle with the given vertices A(1,3), B(-3,5), and C with C = 2A.

(b) Use (a) to find the coordinates of the point D such that the quadrilateral ABCD is a parallelogram?


1
Expert's answer
2022-01-10T15:04:13-0500

a) Let us find the area "A" of the triangle with the given vertices "A(1,3), B(-3,5)," and "C" with "C = 2A=(2,6)."

Taking into account that the equation of the line "AC" is "y=3x" or "y-3x=0," we conclude that the distance (altitude) "h" from the point "B" to the line "AC" is equal to "h=\frac{|5-3(-3)|}{\sqrt{1^2+3^2}}=\frac{14}{\sqrt{10}}." Since "AC=\sqrt{1^2+3^2}=\sqrt{10}," we conclude that

"A=\frac{1}2AC\cdot h=\frac{1}2\sqrt{10}\frac{14}{\sqrt{10}}=7" (sq. units).


b) Let us find the coordinates of the point "D(a,b)" such that the quadrilateral "ABCD" is a parallelogram. Let "O(x,y)" is the point of the diagonals intersection. Then "x=\frac{1+2}2=\frac{3}2,\ y=\frac{3+6}2=\frac{9}2." It follows that "\frac{a-3}2=\frac{3}2,\ \frac{b+5}2=\frac{9}2." Therefore, "a=6,\ b=4."

We conclude that "D(6,4)."